News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 17 August 2001 2.30pm EDT

Last week the observing schedule was interrupted following a high radiation event at 10:08pm EDT on August 15 that stopped the daily load and activated the science instrument safing sequence (SCS 107).

All actions were nominal and assessment of the dump data confirmed that the trigger was due to high proton counts in the EPHIN E1300 channel exceeding the pre-defined threshold of 10 counts/cm^2/s/sr. The solar levels subsided to an acceptable level over the next 24 hours and the loads were replanned and resumed at 8.30pm EDT on 16 Aug. The observations of Cal 83 and UV Seti were impacted and will be rescheduled in a future load. The load was resumed in time to ensure the proper commanding for the first eclipse of the Fall 2001 season at 2.50pm on August 17, and for the observation of a Target Of Opportunity, NGC 6640, that was accepted for observation on August 18.

Real-time procedures were completed on 13 August to load the B-side OTG software patch and to perform a routine update the on-board ephemeris. Procedures were also completed on 17 August to re-enable SCS 107 and re-position the HRC door following the radiation safing action.

Of note last week was the release of an image of Cen-A featured in the Science Section of the August 14 New York Times, and the conduct of a Safe Mode Recovery training simulation on August 17.

The planned observations for the coming week are shown below and include an observation of IC4329A coordinated with RXTE.

------------------------------------------
HD 152404               ACIS-I       Aug 19
HD150193                ACIS-I
PSR J1740+1000          ACIS-S
Vega                    HRC-I
Vega (3 obs)            HRC-S
LBQS 0021-0213          ACIS-S
Radiation Belts                      Aug 20
NGC7619/NGC7626         ACIS-I
3C 452                  ACIS-S       Aug 21
M15                     HRC-I        Aug 22
WZ SGE                  ACIS-S
3C 334                  ACIS-S
Radiation Belts
IC1262                  ACIS-S       Aug 23
A2125                   ACIS-I
3C351                   ACIS-S       Aug 24
Radiation Belts                      Aug 25
IC 4329A                ACIS-S       Aug 26

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports